OpenStudy (anonymous):

The circle formed by the blades of a fan has a circumference of thirty-seven and five-sevenths inches. What is its diameter? Use twenty-two sevenths for pi.

OpenStudy (anonymous):

Circumference = 2(pi)(diameter) so diameter = circumference / 2pi = 264/7 / 2(22/7) = 6 inches